Dear sir it say measons combine with protons and form neutrons please explain the mechanism
1
Expert's answer
2020-10-07T10:13:30-0400

To understand how this is possible, you need to learn about quarks. By now, we know that some particles consist out of quarks. Some - because there are indeed elementary particles (electrons, neutrinos, photons, quarks and gauge bosons). But particles like neutrons, protons and mesons are compound particles. There are 6 quarks - up (u), down (d), strange (s), charm (c), beauty (b) and top(t), and corresponding to them antiquarks.

Quark composition of proton: p=uud; neutron: n=udd. Mesons are particles that consists of only 2 quarks. The lightest mesons are pions.

As an example, we can analyse reaction

"\pi^- + p \rightarrow \pi^0 + n"

"(\bar u d) + (uud) \rightarrow (u\bar u -d \bar d) + (udd)"

As every particle-antiparticle pair, quarks can annihilate, and as opposite quark-antiquark pairs can be born when there is enough energy for creation of its bounded state. Neutral pion is a mixture of two states. As you can see, quark composition remain the same except pair "d \bar d" that appears because the total energy of "\pi^- + p" is enough to create it and, therefore, enough to born heavier particles "\pi^0 + n".